an unknown gas at 51.1 Celsius and 1.10 ATM has a molar mass of 28.01 g/mol assuming ideal behavior what is the density of the gas?
Answer #1

P = 1.1atm

T= 51.1 oC

= (51.1+273) K

= 324.1 K

Lets derive the equation to be used

use:

p*V=n*R*T

p*V=(mass/molar mass)*R*T

p*molar mass=(mass/V)*R*T

p*molar mass=density*R*T

Put Values:

1.1 atm *28.01 g/mol = density * 0.08206 atm.L/mol.K *324.1 K

density = 1.1585 g/L

Answer: 1.16 g/L